### v1

@@ 7 Conclusions
In this paper, we propose an HSI denoising and destriping method, which has an optimization model given in (1) and an iterative algorithm described in Algorithm 1. The optimization model consists of a data fidelity term, a sparsity-enhanced nonlocal low-rank tensor regularization term for denoising, and a $\ell_{2,p}$ norm for destriping. The iterative algorithm is proposed using a proximal version of BCD algorithms, which has convergence guarantees. The numerical experiments tested on simulated and real HSIs also show the effectiveness of our proposed method in removing Gaussian noise, stripes, and deadlines.

### v2

@@ 7 Conclusions
In this paper, we propose an HSI denoising and destriping method, which has an optimization model given in (1) and an iterative algorithm described in Algorithm 1. The optimization model consists of a data fidelity term, a sparsity-enhanced nonlocal low-rank tensor regularization term for denoising, and a $\ell_{2,p}$ norm for destriping. The iterative algorithm is proposed using a proximal version of BCD algorithms, which has convergence guarantees. The proposed model can be extended to a multi-scale low-rank regularized model and the numerical experiments tested on simulated and real HSIs also show the effectiveness of our proposed method in removing Gaussian noise, stripes, and deadlines.